Shaunice has 36 feet of fence to put around her rectangular flowerbed. How long and how wide will the flowerbed have to be so that she has 72 square feet of area to plant flowers?
Answer by mananth(16949)

let length be l
width be w
..
2(l+w)=36
/2
l+w=18
...
Area = l*w = 72
..
12 & 6 are the factors
..
The dimensiona are 12 feet by 6 feet
...
CHECK
Perimeter = 12+12+6+6=36
area = 12*6 = 72
..
m.ananth@hotmail.ca